If two dice are thrown, what is the probability that the first die shows a 4 or that the total on the two dice is 8?
Answer #1

P(First die shows 4 or total is 8)

= P(First die shows 4) + P(Total 8) - P(Both)

= 6/36 + 5/36 - 1/36

= 10/36

= 5/18